OpenStudy (anonymous):

sin^4x+cos^4x=1-2cos^2x+2cos^4x hey can anyone do this trig identity? You'll be awesome :)

OpenStudy (anonymous):

Sin^4x=(sin^2x)^2=(1-cos^2x)^2=1-2cos^2x+cos^4x therfore Left side of eqn is = 1-2co^2x+cos^4x+cos^4x= RHS

OpenStudy (turingtest):

you have the answer, why are you bumping the question?

OpenStudy (anonymous):

maybe he want know how to prove...

OpenStudy (anonymous):

Sorry.. It was a little confusing, but I'll work with I have.

OpenStudy (anonymous):

How did he get 1-cos^2x?

OpenStudy (anonymous):

sin^2(x)= 1-cos^2(x) is fundamental identity of trig